OpenStudy (anonymous):

4.5 qt / min = ____ gal / h a. 67.5 b. 33.8 c. 56.3 d. 65

OpenStudy (the_fizicx99):

xD I think its 67.5 because if there are 4 qts in one gallon then 4.5*60 = 270 / 4 = 67.5

OpenStudy (anonymous):

lol , thanks ! (x

OpenStudy (the_fizicx99):

:3